(b). Use the chain rule to find aw and as y = 8 cost, z = s sint when s= 1 and t=0 aw at where w = = 22 + y2 + z2, x = st,
Answer #1

2-st 0 Giren N W = x+y+zi y = 8 cost 78 Sint ow X ay is + me zex sex me 22 as - me ry OS ar og gw (2x) (+) +(@y)6)+ @2) (0) o